OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9 (1, 12) (−1, −12) (12, −1) (12, 1)

OpenStudy (anonymous):

12, -1

OpenStudy (anonymous):

12, −1)

OpenStudy (anonymous):

Could any of you too by chance show me the work because I'm still not getting it. thanks

OpenStudy (anonymous):

Row reduce this matrix\[ \left[ \begin{array}{ccc} 1 &-2 & 14 \\ 1 & 3 & 9 \\ \end{array} \right]\] to get \[ \left[ \begin{array}{ccc} 1 &0 & 12 \\ 0 & 1 & -1 \\ \end{array} \right]\] Thus x=12 and y=-1 or \[x=\frac{\left|\begin{array}{ccc} 14 & -2 \\ 9 & 3 \\ \end{array} \right|}{\left|\begin{array}{ccc} 1& -2 \\ 1 & 3 \\ \end{array} \right|}=12\] \[y=\frac{\left|\begin{array}{ccc} 1 & 14 \\ 1 & 9\\ \end{array} \right|}{\left|\begin{array}{ccc} 1& -2 \\ 1 & 3 \\ \end{array} \right|}=-1\]

OpenStudy (anonymous):

@cyter:I believe that is also known as cramer's rule ( http://en.wikipedia.org/wiki/Cramer%27s_rule)

OpenStudy (anonymous):

x - 2y = 14................(a) x + 3y = 9.................(b) from (a) x = 14 + 2y...... substitute in b 14 + 2y + 3y =9 5y = -5 y = -1 then substitute in a x - 2(-1) = 14 x = 14 - 2 x = 12 ans: (12, -1)

OpenStudy (anonymous):

Lets use the matrix method: \[ Ax=B \Rightarrow x=A^{1}B \] Here, \[ A = \left[ \begin{array}{ccc} 1 &-2 \\ 1 & 3 \\ \end{array} \right] \] \[ B = \left[ \begin{array}{ccc} 14 \\ 9 \\ \end{array} \right] \] \[ x = \left[ \begin{array}{ccc} x_1 \\ x_2 \\ \end{array} \right] \] Again,\[ A^{-1} = \frac15 \left[ \begin{array}{ccccc} 3 &2 \\ -1 & -1\\ \end{array} \right] \] Giving \[ \left[ \begin{array}{ccc} x_1 \\ x_2 \\ \end{array} \right] =\left[ \begin{array}{ccc} 12 \\ -1 \\ \end{array} \right] \] and we are done here :)

OpenStudy (anonymous):

or graph \[y = \frac{1}{2}x-7\] and \[y = -\frac{1}{3}x+3\] and find the intersection